# Assessing biomedical knowledge robustness in large language models by query-efficient sampling attacks

### v2

@@ 3.3 Sampling in zeroth-order gradient attack
This is an example of the random directions stochastic approximation (Nesterov & Spokoiny, 2017), where $\textbf{u}=(\mathbf{e}_{j}-\mathbf{e}_{0})/\|\mathbf{e}_{j}-\mathbf{e}_{0}\|$ is a randomly oriented vector in the embedding space corresponding to the neighboring entity $e_{j}$. Berger et al. (2021) computed Eq. 6 by random sampling of neighbors (in candidate sampler) within a similarity threshold for synonyms. Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and the discrete nature of the search space, we can rewrite Eq. 5 effectively as deterministic DWS in the embedding space by
@@
4 Implementing attacks on biomedical question answering
@@ 4.1 Dataset selection
We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking. Both datasets are entity-rich and cover a wide range of biomedical specialties and topics. The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States. The MedMCQA dataset contains short questions used in medical exams in India. They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023a; Han et al., 2023; Singhal et al., 2023b). Both datasets are publicly available and don’t contain personal information. We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019). We then divided the QA datasets into drug- or disease-mention questions by the binary mention labels constructed according to the UMLS entity types (see Appendix C).
4.2 Type-consistent entity substitution
Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g. growth hormone, acquired immunodeficiency syndrome).
Algorithm 1 Type-consistent entity substitution (TCES) template for collecting adversarial entities in distractors.
Inputs: Question Q, LLM, entity type $\tau$, query budget $B$.
Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, correct answer k, and evaluator function $\mu$.
Outputs: Number of queries and the replacement entity, None if unsuccessful when the query budget is exhausted.
@@ 4.2 Type-consistent entity substitution
This effectively results in more substitution patterns (one to two words, two to one word, two to two words, two to three words, etc.) than the more common one-to-one synonym substitution (Mozes et al., 2021). We describe type-consistent entity substitution (TCES) in Algorithm 1 as the task-agnostic attack template. TCES is a general and contextualized attack for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, implemented using token boundaries. The template can accommodate single or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
@@ 4.3 Attack execution on LLMs
We selected both generalist (GLMs) and specialist (BLMs) models for robustness evaluation via blackbox adversarial attacks. The criterion we adopted here for a model to have biomedical knowledge is that it should have a baseline performance better than random guessing (e.g. an accuracy of 0.25 for a specific four-choice QA task) in the relevant domain-specific task. For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2024) and UL2 model (Tay et al., 2022) (Flan-UL2). For BLMs, we used MedAlpaca-7B (Han et al., 2023), Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024) and Writer’s Palmyra-Med-20B (Kamble & Alshikh, 2023). The selected models contain diverse encoding schemes and architectures and their sizes are $\sim$ 1-20B.
@@
To single out the effect of the attacks, all models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D). Model inference of Palmyra-Med-20B (Kamble & Alshikh, 2023) used 4-bit quantization to improve speed (Dettmers & Zettlemoyer, 2023). The evaluation settings resulted in our somewhat lower baseline accuracies than the reported ones which are often achieved with the help of few-shot prompting and prompt optimization. The DiscreteZOO attack (Berger et al., 2021) was originally implemented in the textattack framework (Morris et al., 2020b), which we updated to be compatible with span-level perturbation using its boundaries, multi-GPU split-model inference, and decoder-based billion-parameter LLMs by modifying the attack recipe. To demonstrate the generality of the attack settings, we focus on the following three aspects in the benchmarks.
@@
  • Query budget: We used a fixed query budget for three types of model evaluations: (i) Single-query (budget = 1) sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with fixed budgets under 100 per input instance.
@@
  • Text embedding: Both PDWS and the DiscreteZOO attacks require sampling from a text embedding. The counter-fitted GloVe embedding (Mrkšić et al., 2016) was a default for DiscreteZOO, but GloVe is poor for representing specialized vocabulary in the biomedical domain (Wang et al., 2018). Therefore, here we chose two recent embedding models that offer an improved latent representation of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from constrastive learning of general web-scale text data.
@@
  • Hyperparameters: For single-query PDWS attacks, we tuned the primary hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides. For multi-query PDWS attacks, hyperparameter $n$ was briefly tre-uned around its optimal value in the single-query attack. Most tuned hyperparameters fall within [-30, -5] and [5, 30]. Operating the DiscreteZOO attack in a compatible setting for entities requires two key modifications: (i) Use a random word ranking, which allows the attack algorithm to operate in the low-query-budget setting. (ii) Disable the similarity threshold for candidate point selection in gradient estimation because only entity type consistency is required. The ASR was greatly improved after this step. Benchmarks of the DiscreteZOO attacks were carried out with these settings.
@@ 5 Finding and assessing adversarial entities
We conducted failure mode analysis using the attack statistics gathered across models and hyperparameter settings. The following analyses compare methods to identify adversarial entities through query-efficient attacks and uncover their characteristics, including the distribution in the embedding space and their effects on model explainability.
Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq. 3) in model accuracy (top) and diversity index (bottom) of replacement entities in successful attacks. The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) Flan-UL2 on MedMCQA. Disease and drug-mention questions are separated by colors. The vertical dashed line indicates the location of random sampling.
@@
@|l @ P1cm @ |P1.2cm @ P1.2cm @ P2.4cm @ |P1.2cm @ P1.2cm @ P2.4cm |@ MedQA-USMLE MedMCQA
@@
MedAlpaca-7B Drugs 0.474 7.4 9.3 / 10.3 0.433 11.6 13.2 / 12.7
Diseases 0.488 9.9 13.3 / 10.7 0.436 13.5 18.1 / 14.2
@@
Table 1: Single-query (budget = 1) attack success rates (ASRs) on subsets of MedQA-USMLE and MedMCQA datasets by perturbing the drug or disease entity mentions. Zero-shot baseline (Baseline) accuracy is included for reference next to the ASR of sampling attacks using perturbation by random sampling (RandS), powerscaled distance-weighted sampling (PDWS). The “-” and “+” signs indicate the two regimes ($n<0$ and $n>0$). For each model, the highest ASR for a type of entity-mention question in each dataset is bolded.
@@ 5.1 Two regimes of adversarial entities
The embedding space picture of the substitution attack in Fig. 1a provides a geometric interpretation of powerscaled DWS to examine different types of adversarial examples by tuning $n$. We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity. Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$. The diversity index indicates the reusability of adversarial entities which we use as a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a).
The dependence of the ASR and diversity on the hyperparameter $n$ in Fig. 2 indicates two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable. TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.1. Fig. 2 shows that a larger positive $n$ in Eq. 3 samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index. Moreover, there exist local maxima of ASR in both regimes with finite $n$, which were discovered by hyperparameter tuning. We call these behaviors the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information. In both single-query and multi-query attacks, the benchmarks show that the $n<0$ regime of PDWS attacks tends to have a greater influence on model performance than the $n>0$ regime, although prominent counterexamples do exist (MedAlpaca-7B and Llama2-Medtuned-13B (for drug-mention questions), and Flan-T5-xl (for disease-mention questions) in Table 5).
@@
The two-regime effect indicates that the success of the attack may come from two sources. At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure. Whereas at larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a). This explanation may be further supported by examining the diversity index, $\eta_{\text{GS}}$, which shows the dominance of a few replacement entities in successful attacks at large distances. We found these highly reusable adversarial entities by ranking their occurrences in successful attacks by powerscaled DWS. For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)777n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)888technetium Tc 99m exametazime on RxList. For disease names, some of the chromosome deletion syndromes and rare diseases names are more reusable than others. These entities dominate the sampling results in the $n>0$ regime for all LLMs, indicating their transferability and universality.

### v3

@@ 3.3 Sampling view of zeroth-order adversarial attack
This is an example of the random directions stochastic approximation (Nesterov & Spokoiny, 2017), where $\textbf{u}=(\mathbf{e}_{j}-\mathbf{e}_{0})/\|\mathbf{e}_{j}-\mathbf{e}_{0}\|$ is a randomly oriented vector in the embedding space corresponding to the neighboring entity $e_{j}$. Berger et al. (2021) computed Eq. 5 by random sampling of neighbors (in candidate sampler) within a similarity threshold for synonyms. The procedure was iterated once for all ranked attack locations. Given a noisy gradient estimate ^$\widehat{\nabla}_{\mathrm{ZO}}\mathrm{LLM}(\mathcal{Q};e_{0})$ and a discrete search space, we can rewrite Eq. 4 effectively as deterministic DWS in the embedding space by
@@
4 Implementing attacks on biomedical QA
@@ 4.1 Dataset selection
We selected over 9.3k questions from the MedQA-USMLE (Jin et al., 2021) dataset and over 3.8k questions from the MedMCQA (Pal et al., 2022) dataset for benchmarking. Both datasets are entity-rich and cover a wide range of biomedical specialties and topics. The MedQA-USMLE dataset contains long-context questions that are used to assess medical students in the United States. The MedMCQA dataset contains short questions used in medical exams in India. They have been used in recent works for assessing the biomedical knowledge in LLMs (Singhal et al., 2023; Han et al., 2023; Liévin et al., 2024; Saab et al., 2024). Both datasets are publicly available and don’t contain personal information. We annotated the NEs in the two datasets according to the entity types of the Unified Medical Language System (UMLS) (Bodenreider, 2004) using scispaCy (Neumann et al., 2019). We divided the QA datasets into drug- or disease-mention questions according to the UMLS entity types (see Appendix C) of NEs in distractors.
Algorithm 1 TCES attack template for collecting adversarial entities in distractors.
Inputs: Question Q, LLM, entity type $\tau$, budget $B$.
Internals: Number of choices $N_{\text{ch}}$, text t, entity Ent, embedding Emb, key or correct answer k.
Outputs: Number of queries and the replacement entity, None if unsuccessful after all attempts.
@@
4.2 Type-consistent entity substitution (TCES)
Implementing attacks that target entities requires accounting for type consistency and compatibility with simultaneous multi-word (or span-level) operations, because of the inseparable nature of the entity components (e.g. growth hormone, acquired immunodeficiency syndrome). This effectively results in more substitution patterns (one to two words, two to one word, two to two words, two to three words, etc.) than the more common one-to-one synonym substitution (Mozes et al., 2021). We describe TCES in Algorithm 1 as a general and contextualized attack template for entities, with the goal of adversarially modifying model outputs with only one span-level substitution, specified by token boundaries. The template can accommodate single- or multi-query attacks and it is applicable to any task that may be formatted into QA (Gardner et al., 2019).
@@ 4.3 Attack execution on LLMs
We selected both generalist (GLMs) and specialist (BLMs) models for robustness evaluation via blackbox adversarial attacks. The criterion we adopted here for a model to have biomedical knowledge is that it should have a baseline performance better than random guessing (e.g. an accuracy of 0.25 for a specific four-choice QA task) in the evaluated domain-specific task. For GLMs, we used instruction-finetuned T5 (Flan-T5) series of models (Chung et al., 2024) and UL2 model (Tay et al., 2022) (Flan-UL2). For BLMs, we used MedAlpaca-7B (Han et al., 2023), MedLlama-3-8B-v2.0 from John Snow Labs (JSL-MedLM3-8Bv2) [^7] https://huggingface.co/johnsnowlabs/JSL-MedLlama-3-8B-v2.0, Llama2-Medtuned-13B (LM2-Medtuned-13B) (Rohanian et al., 2024), and Palmyra-Med-20B (Kamble & Alshikh, 2023) from Writer. The selected open-source models have sizes in $\sim$ 1-20B range. Besides, we also evaluated GPT-3.5 from OpenAI.
@@
All models were evaluated at zero temperature or in the non-sampling setting and model inference was conducted in the zero-shot setting with only basic prompt instructions (see the prompt structure in Appendix D). Model inference of Palmyra-Med-20B (Kamble & Alshikh, 2023) used 4-bit quantization to improve speed (Dettmers & Zettlemoyer, 2023). The evaluation settings resulted in our somewhat lower baseline accuracies than the reported ones which are often achieved with the help of few-shot prompting and prompt optimization. The DiscreteZOO attack (Berger et al., 2021) was originally implemented in the textattack framework (Morris et al., 2020b), which we updated to be compatible with span-level perturbation using its boundaries, multi-GPU split-model inference, and decoder-based billion-parameter LLMs by modifying the attack recipe. The following three aspects in the benchmarks are in our focus.
@@
  • Query budget: We used fixed query budgets ($B$) for three main types of attacks: (i) Single-query sampling-based attacks were used as the reference because the DiscreteZOO attack requires a minimum of 3 model queries; (ii) Multi-query attacks used a budget of 8 for reasonable computational cost across all models and attack settings for both sampling- and search-based attacks; (iii) The query scaling trends of specific LLMs and different attack settings were investigated with a series of query budgets under 100 per input instance. Here, we also included the deterministic versions of PDWS to sample the nearest and farthest elements with respect to an anchor point. The element was taken from the perturbation set and the distances (between entities) were calculated using cosine in a chosen text embedding. Their query-limited versions are refered to as $B$-nearest and $B$-farthest element sampling, respectively, where all $B$ nearest or farthest elements were used sequentially to attack models in query scaling studies. Detailed definitions are given in Appendix B.4.
@@
  • Text embedding: Both the PDWS and DiscreteZOO attacks require sampling from a text embedding. The counter-fitted GloVe embedding (Mrkšić et al., 2016) was the default for DiscreteZOO, but GloVe is poor for representing biomedical vocabulary (Wang et al., 2018). We chose two recent embedding models with improved latent representations of text relations for the benchmarks: (i) CODER (Yuan et al., 2022), a BERT-structured embedding model contrastively pretrained to distinguish UMLS knowledge graph concepts; (ii) GTE-base (Li et al., 2023b) from contrastive learning of general web-scale text data.
@@
  • Core hyperparameters: For single-query PDWS attacks, we tuned the hyperparameter $n$ within the interval of [-50, 50] using grid search with a step of 5 or 10 because the local maxima of ASR appear on the positive and negative sides. For multi-query PDWS attacks, hyperparameter $n$ was briefly re-tuned around its optimal value in the single-query attack. Most tuned hyperparameters fall within [-30, -5] and [5, 30]. Operating the DiscreteZOO attack in a compatible setting for entities requires two key modifications: (i) Use a random word ranking, which allows the attack algorithm to operate in the low-query-budget setting. (ii) Disable the similarity threshold for candidate point selection in gradient estimation because only entity type consistency is required. The ASR was greatly improved after this step. Benchmarks of the DiscreteZOO attacks were carried out with these settings.
@@ 5 Finding and assessing adversarial entities
Figure 2: Powerscaled DWS of adversarial distractors exhibits a two-regime effect at negative and positive $n$ values (see Eq. 2) in ASR (top) and diversity index (bottom) of replacement entities in successful attacks. Local maxima in ASR with a finite $n$ are also present in each regime. The vertical dashed line indicates the location of random sampling. The observed similar behaviors are compared across models and datasets in (a) Flan-T5-xxl on MedQA-USMLE, (b) Palmyra-Med-20B on MedQA-USMLE, (c) Flan-T5-xl on MedMCQA, (d) MedAlpaca-7B on MedMCQA. Disease and drug-mention questions are separated by colors. The average prompt semantic similarity displayed in (e)-(h) is calculated for the successful attacks obtained from the corresponding attack settings in (a)-(d), respectively.
@@
@|l @ P1.2cm @ |P1.2cm @ P1.2cm @ P2.6cm @ |P1.2cm @ P1.2cm @ P2.6cm |@ MedQA-USMLE MedMCQA
@@
GPT-3.5 (175B) Drugs 0.438 9.1 11.2 / 10.7
Diseases 0.453 8.6 11.5 / 8.8
MedAlpaca-7B Drugs 0.479 7.2 8.2 / 9.6 0.448 13.9 16.2 / 15.9
Diseases 0.498 8.4 12.5 / 9.2 0.434 14.8 19.1 / 13.7
JSL-MedLM3-8Bv2 Drugs 0.827 10.4 10.2 / 11.8
Diseases 0.771 9.5 18.5 / 11.3
@@
Table 1: Single-query (budget = 1) attack success rates (ASRs) on subsets of MedQA-USMLE and MedMCQA datasets by perturbing the drug or disease entity mentions. Zero-shot baseline (Baseline) accuracy is included for reference next to the ASR of sampling attacks using perturbation by random sampling (RandS), powerscaled distance-weighted sampling (PDWS). The “-” and “+” signs indicate the two regimes ($n<0$ and $n>0$). For each model, the highest ASR for a type of entity-mention question in each dataset is bolded. The models with an asterisk () were only evaluated on one dataset to limit the computational cost.
@@ 5.1 Two regimes of adversarial entities
The embedding space picture of the substitution attack in Fig. 1a provides a geometric interpretation of PDWS to examine different types of adversarial examples by tuning $n$. We use the Gini-Simpson index (Rao, 1982), $\eta_{\text{GS}}=1-\sum_{j}q_{j}^{2}$, to quantify the diversity of adversarial examples and a low value indicates less diversity. Here, $q_{j}\in(0,1]$ is the probability of occurrence of the entity $e_{j}$ among all instances of adversarial entities obtained from sampling, and $\sum_{j}q_{j}=1$. A low diversity index indicates the reusability and is a proxy indicator for the universal adversarial attack regime (Moosavi-Dezfooli et al., 2017; Wallace et al., 2019a). Combining the ASR dependence on diversity and hyperparameter $n$ in Fig. 2 reveals two regimes: (i) the adversarial entities semantically close to ($n<0$) the key $k$ have a greater diversity; (ii) the adversarial entities semantically far from ($n>0$) the key $k$ are more reusable. TCES in these two regimes both can lead to an increase in successful attacks from random sampling (equivalent to $n=0$), as shown in Tables 5-5.2. Fig. 2 shows that a larger positive $n$ in Eq. 2 samples more universal adversarial entities than in other situations, leading to a marked drop in the diversity index. Moreover, there exist local maxima of ASR in both regimes with finite $n$, which were discovered by hyperparameter tuning. We call the behavior the two-regime effect and they were observed in almost all models evaluated here in varying magnitudes for both QA datasets, thanks to the controllable spatial coverage of sampling methods incorporating distance information. In both single-query and multi-query attacks, the benchmarks show that the $n<0$ regime of PDWS attacks tends to have a greater influence on model performance than the $n>0$ regime, although prominent counterexamples do exist (MedAlpaca-7B and Llama2-Medtuned-13B (for drug-mention questions), and Flan-T5-xl (for disease-mention questions) in Table 5).
@@
Semantic distortion by adversarial entities
Besides ASR, evaluating the semantic distortion is another way to quantify the effectiveness of adversarial attacks. We estimated the semantic distortion of entire prompts in successful attacks before and after entity substitution, corresponding to $\mathcal{Q}$ and $\mathcal{Q}^{\prime}$ in Definition 3.2. We calculated the average prompt semantic similarity (PSS) using SentenceTransformer (Reimers & Gurevych, 2019) with the RoBERTa-large model (all-roberta-large-v1). At each distinct $n$, the PSS $\in[-1,1]$ with 1 being the most similar, is averaged over all successful attack instances to obtain the average PSS, a quantitative measure of the semantic distortion. Its dependence on the hyperparameter $n$ for different models and datasets are shown in Fig. 2.
@@
The two-regime effect indicates that the success of the attacks may come from two sources that dominate at different locations on the same attack surface, which refers to entity substitution in the present work.
  • At larger distances, the attack succeeds potentially due to the obscurity of the distractor entities in the question context, which can also profoundly impact the model performance (Li et al., 2023a). This explanation may be further supported by examining the diversity index, $\eta_{\text{GS}}$, which shows the dominance of a few replacement entities in successful attacks at large distances. We found these highly reusable adversarial entities by ranking their occurrences in successful attacks by powerscaled DWS. For drug names, the top-ranked adversarial entities include n-acetylglucosamine (an amino sugar and anti-inflammatory drug)888n-acetylglucosamine on RxList and technetium Tc 99m exametazime (a radiopharmaceutical and contrast agent)999technetium Tc 99m exametazime on RxList. For disease names, some of the chromosome deletion syndromes and rare disease names are more reusable than others. These entities dominate the sampling results in the $n>0$ regime for all LLMs, indicating their transferability and universality.
  • At small distances, semantic proximity (Mozes et al., 2021) reduces the distinguishability between the key and the distractor, therefore increasing the chance of model failure. Examples include similar drugs or diseases with insufficient details such as using diabetes as a distractor instead of the correct answer diabetes II (refers to type-II diabetes) can result in a successful attack.
  • Adversarial distractors result in little semantic distortion at the prompt level (see Fig. 2e-h). The experiemnts on the MedMCQA dataset yields somewhat lower average PSS than MedQA-USMLE because of the longer format of the latter dataset, leading to sightly larger distortions by entity substitution. Adversarial entities obtained in the $n<0$ regime results in slightly higher average PSS compared to those in the $n>0$ regime, while there is also a two-regime effect – the average PSS tends to be the lowest around $n=0$ and higher when $n$ departs further from 0. The trend is consistent for different models and datasets demonstrated in Fig. 2.
5.2 Scaling characteristics from adversarial entities
Size is not all for scaling model robustness.
We compare model robustness using single-query ASR as the metric in Table 5 and Fig. 3. For the Flan-T5 series of GLMs (Chung et al., 2024) and Flan-UL2 (Tay et al., 2022), the single-query ASR is more pronounced in smaller and less performant models, indicating the improvement of robustness as the model’s size grows. However, an opposite trend of robustness is seen in BLMs. It is worth noting that although Palmyra-Med-20B (Kamble & Alshikh, 2023) has about three times the number of parameters than MedAlpaca-7B (Han et al., 2023), the larger model is noticeably more sensitive to entity perturbation. In terms of ASR, the BLMs are comparable to the GLMs in each type of questions evaluated on. Another observation is that the ASR for models is asymmetrical under PDWS attacks. For most of the models, sampling adversarial entities with inverse-distance-scaled weights ($n<0$ in Eq. 2) induces more performance drop than sampling with distance-scaled weights ($n>0$ in Eq. 2), yet prominent counterexamples exist. The attack performance is further investigated in the sequential, multi-query setting (budget = 8) in Table 5.2, where we also compared sampling- against search-based DiscreteZOO attacks. Here, PDWS attacks in the $n<0$ regime are still largely the most effective.